# Spinodal decomposition in rapidly expanding quark gluon plasma

### v1

@@ Abstract
Quantum Chromodynamics (QCD) is expected to have a first order phase transition between the confined hadron gas and the deconfined quark gluon plasma at high baryon densities. This will result in phase boundary effects in the metastable and unstable regions. It is important to include these effects in phenomenological models of heavy ion collisions to identify experimental signatures of a phase transition. This requires building intuition on phase separation in rapidly expanding fluids. In this work we present the equations of relativistic hydrodynamics with a phase boundary, provide prescriptions to extend the equation of state to metastable and unstable regions, and show the effects of spinodal separation in a Bjorken flow.
@@ I Introduction
Simulations have shown that even individual BES collisions span a wide region within the phase diagram and the volume of matter whose trajectory is in the neighbourhood of the critical point is not necessarily very large [38, 39, 40]. This makes it even more challenging to get clean experimental signature of such a critical region. What may be more realistic to achieve is to detect effects of phase boundary created by a first order phase transition. As this is a wide region on the phase diagram, it is much more plausible that a collision of suitable energy will have a big chunk of its volume passing through the phase transition line. Of course this brings in its own challenges. The phase transition can proceed by two paths: nucleation or spinodal decomposition. Once the system supercools nucleation is the first process to initiate the phase change [41]. Generally this will happen in the metastable region. However, if the nucleation rate is small relative to the expansion rate then the system will enter the unstable region and spinodal decomposition will occur [42]. Spinodal decomposition may be initiated by thermal fluctuations or by fluctuations already present in the system. Most simulations assume the latter and proceed to solve the Cahn-Hilliard equation. Obviously there are initial state fluctuations present in heavy ion collisions. While the first order phase transition has been extensively studied in chemical and condensed matter systems (for example, see the textbook [43]), the relativistic collision systems bring in an additional problem in that it is rapidly expanding. To the best of our knowledge, the effects of phase separation in a rapidly expanding fluid have not been explored.
As a first step towards this goal, following the pioneering work of Ref. [44, 45], we write the equations of spinodal decomposition in a rapidly expanding fluid. We provide a prescription to extend the EOS to the metastable and unstable regions, and show the results of numerical simulations of a 1D rapidly expanding QCD medium undergoing the first order phase transition. The thermodynamics of first order phase transition is discussed in Sec. II. We provide prescriptions to extend the equation of state to metastable regions in Sec. III and write the equations of hydrodynamics with a phase boundary in Sec. IV. The equations for the case of Bjorken flow are given in Sec. V and we show the effects of first order phase transition on Bjorken fluid in Sec. VI. We discuss our findings in Sec. VII.
@@ V Hydrodynamics: Bjorken
In boost-invariant hydrodynamics one expresses the time and location along the beam direction in terms of the proper time $\tau$ and space-time rapidity $\xi$ as
@@
Since diffusion is probably necessary for spinodal decomposition, or phase separation, we should allow the thermodynamic functions to depend on $\xi$ as well as $\tau$. The energy-momentum conservation equation is
$\partial_{\mu}T^{\mu\nu}$ $=$ $\left(\frac{\partial\tilde{w}}{\partial\tau}+\frac{\tilde{w}}{\tau}\right)u^{\nu}-g^{\mu\nu}\partial_{\mu}\tilde{P}+K\left[D^{\nu}(\partial_{\mu}D^{\mu}n)+D^{\mu}(\partial_{\mu}D^{\nu}n)\right]=0.$ (36)
In what follows we allow the baryon density to depend on $\xi$ but the energy density and pressure are taken to be boost invariant. As the system evolves, the baryon density variation will break the boost invariance of the energy density and pressure which would lead to an acceleration term. In this work we treat this as a perturbation and neglect the feedback. As we shall see, this is justified in our numerical example. The implications of a fully boost-invariace breaking general system of equations is left for future work. Some useful expressions are
@@ VI Spinodal separation in Bjorken hydrodynamics
We initialize the system at time $\tau_{x}=10$ fm such that the energy density $\epsilon(\tau_{x})$ is $\xi$-independent while the density has the periodic fluctuation $n_{fl}\sin(k\xi)$ around the average value $n(\tau_{x})$. The fluctuation is necessary to seed spinodal decomposition. These are chosen such that the system just reaches the coexistence curve. Specifically, the point $\epsilon(\tau_{x})$, $n(\tau_{x})-n_{fl}$ lies on the coexistence curve. We evolve the system following Eqs. (48-50).
@@ VII Discussion
In this manuscript we have incorporated terms into the stress-energy-momentum tensor, which are normally included to describe spinodal decomposition in non-relativistic systems, into covariant hydrodynamics. These terms are second-order in thermodynamic densities, namely, the baryon density and its space-time derivatives. Only one new parameter, $K$, enters and is associated with the surface free energy between the two phases. It is also necessary to have a parameterization of the equation of state when the system is in the metastable or unstable regions. This is better accomplished by expanding in powers of the baryon density rather than the volume per baryon. The resulting equations of motion are fourth order in space-time derivatives. This makes solution of the space-time evolution very numerically challenging. Therefore, in this manuscript we only gave a simple illustration of how this works in a nearly boost-invariant Bjorken scenario. Future work should include realistic initial state and hydrodynamic fluctuations. Only then will comparisons of theoretical simulations with experimental observations fulfill our goal. Given advances in computational calculations over the past two decades we are optimistic that the goals can be achieved.

### v2

# Covariant formulation of spinodal decomposition in rapidly expanding quark gluon plasma

@@ Abstract
Quantum Chromodynamics (QCD) is expected to have a first order phase transition between the confined hadron gas and the deconfined quark gluon plasma at high baryon densities. This will result in phase boundary effects in the metastable and unstable regions. It is important to include these effects in phenomenological models of heavy ion collisions to identify experimental signatures of a phase transition. This requires building intuition on phase separation in rapidly expanding fluids. In this work we present the covariant equations of relativistic hydrodynamics with a phase boundary, provide prescriptions to extend the equation of state to metastable and unstable regions, and show the effects of spinodal separation in a Bjorken flow.
@@ I Introduction
Simulations have shown that even individual BES collisions span a wide region within the phase diagram and the volume of matter whose trajectory is in the neighbourhood of the critical point is not necessarily very large [38, 39, 40]. This makes it even more challenging to get clean experimental signature of such a critical region. What may be more realistic to achieve is to detect effects of phase boundary created by a first order phase transition. As this is a wide region on the phase diagram, it is much more plausible that a collision of suitable energy will have a big chunk of its volume passing through the phase transition line. Of course this brings in its own challenges. The phase transition can proceed by two paths: nucleation or spinodal decomposition. Once the system supercools nucleation is the first process to initiate the phase change [41]. Generally this will happen in the metastable region. However, if the nucleation rate is small relative to the expansion rate then the system will enter the unstable region and spinodal decomposition will occur [42]. Spinodal decomposition may be initiated by thermal fluctuations or by fluctuations already present in the system. Most simulations assume the latter and proceed to solve the Cahn-Hilliard equation. Obviously there are initial state fluctuations present in heavy ion collisions. While the first order phase transition has been extensively studied in chemical and condensed matter systems (for example, see the textbook [43]), and has been explored in the non-relativistic limit of low energy nuclear collisions [44, 45], the relativistic collision systems bring in an additional problem in that it is rapidly expanding to relativistic velocities. There has been some pioneering work on exploring the spinodal region in nuclear collisions [46, 47, 48, 49, 50, 51, 52] but, to the best of our knowledge, a covariant energy-momentum tensor including the surface energy has not been written. This is also the first work exploring the effects of phase separation in a fluid rapidly expanding in the longitudinal direction.
As a first step towards this goal, we write the equations of spinodal decomposition in a rapidly expanding fluid. We provide a prescription to extend the EOS to the metastable and unstable regions, and show the results of numerical simulations of a 1D rapidly expanding QCD medium undergoing the first order phase transition. The thermodynamics of first order phase transition is discussed in Sec. II. We provide prescriptions to extend the equation of state to metastable regions in Sec. III and write the equations of hydrodynamics with a phase boundary in Sec. IV. The equations for the case of Bjorken flow are given in Sec. V and we show the effects of first order phase transition on Bjorken fluid in Sec. VI. We discuss our findings in Sec. VII.
@@ V Hydrodynamics: Bjorken
In boost-invariant hydrodynamics [56, 57, 58] one expresses the time and location along the beam direction in terms of the proper time $\tau$ and space-time rapidity $\xi$ as
@@
Diffusion is necessary for spinodal decomposition, or phase separation, as it facilitates the flux from lower to higher density region when the chemical potential gradient is opposite to the density gradient in the unstable region. So we should allow the thermodynamic functions to depend on $\xi$ as well as $\tau$. The energy-momentum conservation equation is
$\partial_{\mu}T^{\mu\nu}$ $=$ $\left(\frac{\partial\tilde{w}}{\partial\tau}+\frac{\tilde{w}}{\tau}\right)u^{\nu}-g^{\mu\nu}\partial_{\mu}\tilde{P}+K\left[(D^{\nu}n)(\partial_{\mu}D^{\mu}n)+(D^{\mu}n)(\partial_{\mu}D^{\nu}n)\right]=0.$ (36)
In what follows we allow the baryon density to depend on $\xi$ but the energy density and pressure are taken to be boost invariant. As the system evolves, the baryon density variation will break the boost invariance of the energy density and pressure which would lead to an acceleration term. In this work we treat this as a perturbation and neglect the feedback. As we shall see, this is justified in our numerical example. The implications of a fully boost-invariance breaking general system of equations is left for future work. Some useful expressions are
@@ VI Spinodal separation in Bjorken hydrodynamics
We initialize the system at time $\tau_{x}=10$ fm such that the energy density $\epsilon(\tau_{x})$ is $\xi$-independent while the baryon density has the periodic fluctuation $n_{fl}\sin(k\xi)$ around the average value $n(\tau_{x})$. This is obviously a simplification. However, it is motivated by the fact that while the energy density is boost invariant around mid-rapidity in most initial state models, the baryon density is not [40, 63]. The fluctuation is necessary to seed spinodal decomposition. These are chosen such that the system just reaches the coexistence curve. Specifically, the point $\epsilon(\tau_{x})$, $n(\tau_{x})-n_{fl}$ lies on the coexistence curve. We evolve the system following Eqs. (48-50).
@@ VII Discussion
In this manuscript we have incorporated terms into the stress-energy-momentum tensor, which are normally included to describe spinodal decomposition in non-relativistic systems, into covariant hydrodynamics. These terms are second-order in thermodynamic densities, namely, the baryon density and its space-time derivatives. Only one new parameter, $K$, enters and is associated with the surface free energy between the two phases. It is also necessary to have a parameterization of the equation of state when the system is in the metastable or unstable regions. This is better accomplished by expanding in powers of the baryon density rather than the volume per baryon. The resulting equations of motion are fourth order in space-time derivatives. This makes solution of the space-time evolution very numerically challenging. Therefore, in this manuscript we only gave a simple illustration of how this works in a nearly boost-invariant Bjorken scenario. Motivated by various initial state models, we have used the simplification that the energy density is boost-invariant while the baryon density fluctuates. Once the system is initialized this way, the violation of boost invariance of the energy density is less than 3% for the system considered here. The pressure gradient terms are ignored. If they were included, the violation of boost invariance would likely be even smaller as those terms would further equilibrate the system. A small value of K was chosen to avoid breaking the smoothly varying nature of the baryon density. Future work should include realistic initial state, hydrodynamic fluctuations, and any other relevant terms. Only then will comparisons of theoretical simulations with experimental observations fulfill our goal. Given advances in computational calculations over the past two decades we are optimistic that the goals can be achieved.